# Evaluating and Mitigating Number Hallucinations in Large Vision-Language Models: A Consistency Perspective

### v1

@@ 4.3 Results and Analysis
We train with different consistency formats to fully reveal the performance of consistency method. Consistency(I) outperforms Direct on some metrics, which has already shown the effectiveness of consistency method. (For detailed discussions, please refer to Appendix D). However, Consistency(II) does not perform as well, which we believe is because comparison only is too difficult and far from our primal task, making it less effective when tested with primal task.
@@
Further, after combining prompt I and II, Consistency(I+II) outperforms Direct on almost all metrics and also beats Consistency(I) and Consistency(II) on most metrics, which highlights the effectiveness of our consistency method and the importance of both perspectives. Although either perspective only does not offer good enough results, combining them leads to an excellent result. This observation also meets our assumption that adding more perspectives may further ease the inconsistency problems and resulting in better overall performance. Our method also bears several advantages such as focusing on task and data only, no need for finetuning the LLM part of LVLMs, etc. For details about the advantages of our method and a case study, please refer to Appendix D.
5 Related Works
5.1 Large Vision-Language Models
With the success of large language models (LLMs), there has been a surge in efforts to incorporate additional modalities into LLMs, leveraging their powerful recognition capabilities. BLIP-2 (Li et al., 2023a), for instance, introduces the Q-former to align image input into tokens suitable for utilization by LLMs. LLaVA (Liu et al., 2023b) and LLaVA-v1.5 (Liu et al., 2023a) employ visual instruction tuning and leverage simple linear layers or multi-layer perceptrons to map image output into the text feature space. InstructBLIP (Dai et al., 2023) achieves instruction following abilities by performing instruction tuning on the BLIP-2 backbone. Notably, MiniGPT-4 (Zhu et al., 2023) and MiniGPT-v2 (Chen et al., 2023a) are also powerful LVLMs that exhibit strong performance across various vision-language tasks. Among all open-source and closed-source LVLMs, GPT-4V (OpenAI et al., 2023) retains its status as the most powerful one.
Hallucination is an extensively discussed phenomenon in both traditional vision-language models and large language models and numerous studies have delved into the examination of hallucinations in large vision-language models (LVLMs). The most extensively discussed form of hallucination is object hallucination, centering around the determination of whether a specific object exists in a given image. In the context of caption-based evaluation, CHAIR (Rohrbach et al., 2018) stands out as a classic metric that compares objects mentioned in the caption with those present in the provided image. POPE (Li et al., 2023b), on the other hand, employs a binary classification approach by directly asking the model whether a particular object exists in the provided image. Zhou et al. (2023) conducted an analysis of factors influencing object hallucination and introduced LURE based on their findings. Leng et al. (2023) proposed visual contrastive decoding to mitigate object hallucinations, while Zhai et al. (2023) suggested a switch-based method to address this issue. However, these studies predominantly focus on the mere existence of objects, which may not be sufficient for gauging whether the model truly comprehends the image.
Beyond object hallucination, there are also studies addressing more generalized forms of hallucinations. Wang et al. (2023a) introduced LLMs to aid in the detection and evaluation of hallucinations in captions. Gunjal et al. (Gunjal et al. (2023)) proposed a dataset and reward model to assess and mitigate hallucinations in captions. RAH (Chen et al., 2023b) and FGHE (Wang et al., 2023b) are two datasets similar to POPE that can be employed for evaluating fine-grained hallucinations. However, many caption-based hallucination evaluation methods require a lot of training modules, so the performance and stability of them remain uncertain due to the complexity of captions and training procedure. POPE-like evaluation metrics rely on binary classification questions and might not fully capture the model’s ability, making them insufficient for quantitative comparisons.
5.2 Counting Task in VQA
There have been studies specifically addressing the task of counting objects in visual question-answering. Earlier works such as Lempitsky and Zisserman (2010) and Segui et al. (2015) laid some groundwork in this area. Zhang et al. (2018) proposed a counting module that can be integrated into other visual question-answering models. Additionally, Liu et al. (2020) introduced a blockwise counting method to tackle issues like inaccurate generated regression targets and imbalanced samples, which were significant challenges in previous counting methods. However, these methods often require specific designs for the task and modifications to the entire model architecture, which may be insufficient and inefficient in the context of large models nowadays.
6 Conclusion
@@
7 Limitations
There are several limitations in our work. First of all, we only use one dataset for discussion instead of using multiple datasets. Secondly, under the constraint of cost we do not perform a quantitative evaluation on GPT-4V, which is a pity. Lastly, we mainly focus on VQA tasks and ignore captioning tasks for its complexity, while we do believe that captioning tasks are still worth discussing.
8 Ethics Statement
We do not perform any specified selection on data or models. All parts of our work are done evenly and automatically. And since we use open-source dataset (MSCOCO) and model weights (LVLMs) just consistently with their intended use (object detection and language generation), we believe there are no ethical problems in our work. ChatGPT is used only to assist writing (grammar checking) in this paper, which we believe is totally acceptable.
Appendix A Details About Dataset Construction
When constructing dataset, we only retain data $(i,c,n_{c})$ with $n_{c}>0$. There are two main considerations for this choice. First of all, there are many objects that do not appear in a given image and retaining all these questions is too expensive. If we perform k-uniform sampling similarly, this may result in randoml selection, which is insufficient for reflecting the model’s performance according to (Li et al., 2023b). More importantly, we conduct an experiment on LLaVA with randomly selected $(i,c,n_{c})$ with $n_{c}=0$ and find out that LLaVA correctly answers 95% of these questions, which meets the its high POPE evaluation result (Liu et al., 2023a). Based on these considerations, we only retain data $(i,c,n_{c})$ with $n_{c}>0$ to focus on the number of an existing object.
We present double k-uniform sampling in algorithm 1. Adjusting the value of $k$ allows us to strike a balance between achieving a more uniform distribution and retaining a larger dataset. In the subsequent discussion, we opt for $k=50$, resulting in a total of 20,200 data points. The distribution of our dataset is shown in Table 6.
Algorithm 1 Double k-Uniform Sampling
1:  Input dataset $S=\{(i,c,n_{c})\}$ and $k$.
2:  Randomly shuffle the dataset $S$.
3:  Initialize $m(c,n)\leftarrow 0,S^{{}^{\prime}}=\phi$
4:  for $j=1,2...,|S|$ do
5:     if $m(c_{j},n_{cj})>k$ then
6:        continue
7:     end if
8:     $S^{{}^{\prime}}\leftarrow S^{{}^{\prime}}\cup\{(i_{j},c_{j},n_{cj})\}$
9:     $m(c_{j},n_{cj})\leftarrow m(c_{j},n_{cj})+1$
10:  end for
11:  Output $S^{{}^{\prime}}$
Answer Percentage
1 19.8%
2 18.7%
3 15.1%
4 11.9%
5 8.0%
6 6.1%
7 4.8%
8 3.7%
9 3.1%
$\geq$ 10 8.8%
Table 6: The distribution of the answers of the questions
As can be seen from the Table 6, though k-uniform sampling is performed, questions with answer $\geq 10$ is still too less. Therefore to avoid randomness, when calculating F-scores, questions with answers $\geq 10$ are considered one category.
Appendix B Inconsistency Examples
As is mentioned, inconsistency is a common problem. Here are some typical responses containing inconsistency.
Prompt: There are 2 motorcycles in this picture, is that correct? Answer yes or no. Prompt: There is 1 motorcycle in this picture, is that correct? Answer yes or no.
Answer: Yes Answer: Yes
Table 7: Example of model answering ’yes’ to contradictory questions.
Prompt: Which object has a larger number in this picture, backpack or dog? If they have the same number, answer same. Prompt: Which object has a larger number in this picture, dog or backpack? If they have the same number, answer same.
Answer: Dog Answer: Backpack
Table 8: Example of model’s respond changing after the order of objects changes.
Prompt: Please select the correct option according to the given picture. A. There are more dogs than bicycles in this picture. B. There are more bicycles than dogs in this picture. C. The number of dogs and bicycles are the same in this picture. Prompt: Please select the correct option according to the given picture. A. There are more bicycles than dogs in this picture. B. There are more dogs than bicycles in this picture. C. The number of bicycles and dogs are the same in this picture.
Answer: B Answer: B
Table 9: Example of model’s response changing after the order of objects changes.
Particularly, Table 9 are two responses generated by MiniGPT-v2. As can be seen, they bear the same input image and prompt with similar meanings and the only difference between them is the order of the objects (dog and bicycle). However, model responds with contradictory answers. In the first case, model correctly identifies there being more bicycles than dogs. But in the second case, the model continues to answer ’B’, which means it believes there are more dogs than bicycles. This inconsistency is a giant problem because if we only see the first case, we may happily assume that the model correctly understands this picture, while only with both cases can we observe this dilemma.
Appendix C Dataset, Models and Hyper-Parameters of Finetuning
Dataset
We obtain training set $S$ in the same way as mentioned in Section 2.2. This provides us with 233k pieces of data. For Direct method, all these data are used by directly asking like Section 2.1.
For Consistency(I), all these data are kept with prompt I, which provides us with 233k pieces of data similarly. As for Consistency(II), 100k pieces of data are randomly selected with prompt II. By combining these two datasets, we get a dataset with 333k pieces of data, corresponding to Consistency(I+II). Note that all these consistency data are automatically constructed from the whole original MSCOCO2014 training set.
Models
For LLaVA, the alignment part is a two-layer MLP. For MiniGPT-v2, the alignment part is a linear layer. For InstrcutBLIP, though the whole alignment part is a Q-former, we perform an experiment and observe that only finetuning the linear layer after Q-former instead of the whole Q-former offers a better performance on the direct training method, so we also only finetune the linear layer after Q-former in InstructBLIP.
Hyper-Parameter
@@ Appendix C Dataset, Models and Hyper-Parameters of Finetuning
Table 10: Main hyper-parameters used for tuning LLaVA.
@@
Table 11: Main hyper-parameters used for tuning InstructBLIP.
@@
Table 12: Main hyper-parameters used for tuning MiniGPT-v2
Appendix D Discussion about Our Method
Discussions
We observe an interesting phenomenon that Consistency+I often offer better weighted-F1 and worse MAE, while Consistency(I+II) may bear a littlt bit worse weighted-F1 with much better MAE than Consitency(I). We believe this phenomenon is because of the distribution of our training set. Since we do not perform double k-uniform sampling on our training set as mentioned, there are much more data with small answers rather than large answers, making the model more confident about generating smaller answers after introducing consistency with binary classification questions. However, when comparison question is introduced, the model is forced to focus on the whole image instead of trying to bypass by generating smaller answers, which result in a small drop in weighted-F1 while an increase in macro-F1 and MAE. This observation also highlights the importance of using these three metrics instead of one particular metric. As can be seen, using three metrics offers a more comprehensive understanding of the model’s performance.
Advantages of Our Method
Case Study
Answer of Direct: 3. Answer of Consistency(I+II): 5.
Table 13: A successful example of consistency method. The gold answer is $5$.
Table 13 shows a successful example of our consistency method. This is also a case that we use to illustrate inconsistency problem in Appendix B, which further prove the effectiveness of introducing consistency into training. As can be seen, with Direct method the model fails to identify the correct number of bicycles, while after introducing Consistency(I+II) method the model responds the correct answer. We believe this improvement is led by our consistency method forces the model to think more carefully and comprehensively, and thus reduces uncertainty and confusion.

### v4

@@ 5.3. Results and Analysis
We train with different consistency formats to fully reveal the performance of consistency method. Consistency(I) outperforms Direct on some metrics, which has already shown the effectiveness of consistency method. However, Consistency(II) does not perform as well, which we believe is because comparison only is too difficult and far from our primal task, making it less effective when tested with primal task.
We observe an interesting phenomenon that Consistency(I) often offer better weighted-F1 and worse MAE as well as macro-F1, while Consistency(I+II) may bear a little bit worse weighted-F1 with much better MAE than Consistency(I). We believe this phenomenon is because of the distribution of our training set. Since we do not perform double k-uniform sampling on our training set as mentioned, there are much more data with small answers rather than large answers, making the model more confident about generating smaller answers after introducing consistency with binary classification questions. But this actually harms the overall performance because it makes larger mistakes when counting more objects. However, when comparison question is introduced, the model is forced to focus on the whole image instead of trying to bypass by generating smaller answers, which result in a small drop in weighted-F1 while an increase in macro-F1 and MAE. This observation also highlights the importance of using these three metrics instead of one particular metric. As can be seen, using three metrics offers a more comprehensive understanding of the model’s performance.
@@
Further, after combining prompt I and II, Consistency(I+II) outperforms Direct on almost all metrics and also beats Consistency(I) and Consistency(II) on most metrics, which highlights the effectiveness of our consistency method and the importance of both perspectives. Although either perspective only does not offer good enough results, combining them leads to an excellent result. This observation also meets our assumption that adding more perspectives may further ease the inconsistency problems and resulting in better overall performance.
Advantages
Case Study
Answer of Direct: 3. Answer of Consistency(I+II): 5.
Table 8. A successful example of consistency method. The gold answer is $5$. This example is drawn from InstructBLIP.
Table 8 shows a successful example of our consistency method. This is also a case that we use to illustrate inconsistency problem in 4.2, which further proves the effectiveness of introducing consistency into training. As can be seen, with Direct method the model fails to identify the correct number of bicycles, while after introducing Consistency(I+II) method the model responds the correct answer. We believe this improvement is led by our consistency method forces the model to think more carefully and comprehensively, and thus reduces uncertainty and confusion.
6. Conclusion
@@
Appendix A Analysis of GPT-4V
Since our test set is rather large, we randomly select 100 pieces of data from our test set to perform a quantitative evaluation of number hallucination on GPT-4V. The evaluation results are listed in Table 9.
Model Macro-F1 Weighted-F1 MAE
Random 0.086 0.103 3.56
LLaVA 0.186 0.288 1.84
LLaVA-L 0.196 0.299 1.80
InstructBLIP 0.201 0.300 2.35
MiniGPT-v2 0.193 0.293 2.06
GPT-4V 0.276 0.447 1.17
Table 9. Number hallucination evaluation results of different LVLMs. GPT-4V is evaluated on only 100 pieces of data.
As can be seen from Table 9, GPT-4V outperforms other models a lot on all metrics, which meets our expectation that GPT-4V is more powerful and bears less hallucination. However, GPT-4V still suffers from number hallucination, and an example is shown in Figure 2.
Figure 2. An example of number hallucination in GPT-4V.
The prompt is ‘How many cars are there in this picture? Answer in a single number.’ It is obvious that there are only 2 cars in this picture, however GPT-4V responds with 4. This example and the quantitative evaluation results clearly demonstrate that even an LVLM as powerful as GPT-4V still suffers from number hallucination, which further emphasizes the importance of discussing and addressing this problem.
Appendix B Hyper-Parameters of Finetuning
@@
Table 10. Main hyper-parameters used for tuning LLaVA.
@@
Table 11. Main hyper-parameters used for tuning InstructBLIP.
@@
Table 12. Main hyper-parameters used for tuning MiniGPT-v2